• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name A. Camsonne

    Log entry time 23:06:41 on February 9,2009

    Entry number 260122

    keyword=Compton photon rate odd

    I was called in to check on the compton which was crashing.
    The compton seemed to work fine after a reboot of coda though I noticed some rate on the central crystal even without beam.
    There seem to be some noise on the central crystal so for now I don't think the photon rate is reliable. I am playing with the gain, attenuation and threshold to try to see if I am able to see signal from the detector. I will call MCC if I manage to get a reliable rate